The demand function for product X is: Qd = 600 – 20Px + 0.02Y – 5Pr The supply function is: Qs = –300 + 10Px
Where:
Qd = the quantity of X demanded Qs = the quantity of X supplied Px = the price of product X Y = the average consumer income Pr = the price of the related product
1
Expert's answer
2019-03-12T10:54:46-0400

Qd = 600 – 20Px + 0.02Y – 5Pr,

Qs = –300 + 10Px.

In equilibrium Qd = Qs, so:

600 – 20Px + 0.02Y – 5Pr = –300 + 10Px,

30Px = 900 - 5Pr + 0.02Y,

Px = 30 - 1/6Pr + 0.01/15Y,

Qx = -300 + 300 - 5/3Pr + 0.1/15Y.